OpenStudy (anonymous):

integrate with respect to x. [x^3+4x^2] dx b=0 a=-4

OpenStudy (anonymous):

is this a definite integral from -4 to 0, then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright, so with any definite integral, first you have to find the indefinite integral (or the antiderivative)

OpenStudy (anonymous):

this would be: \[x^4/4 + 4x^3/3\]

OpenStudy (anonymous):

now you plug in the higher value of x into the equation (you get 0)

OpenStudy (anonymous):

you the subtract the indefinite integral evaluated at the lower value of x. this happens to be: 64-256/3

OpenStudy (anonymous):

so your final result is 256/3 - 64

OpenStudy (anonymous):

remember that you subtracted

OpenStudy (australopithecus):

Split the integral \[\int\limits_{-4}^{0} x^{3} +\int\limits_{-4}^{0} 4x^{2}\] you will get: \[x^{4}/4 + 4x^{3}/3 \] then use b - a so \[\frac{0^{4}}{4} - \frac{40^{3}}{3} - (\frac{(-4)^{4}}{4} - \frac{4(-4)^{3}}{3})\]

OpenStudy (australopithecus):

compute that for your answer

OpenStudy (anonymous):

I already did :P

OpenStudy (anonymous):

but I would say that your answer is more formal

OpenStudy (anonymous):

you may have wanted to put a parenthesis around the 0 in the second term, though. (its not 40^3)